A bag of peanuts is worth $0.20 less than a bag of cashews. Equal amounts of peanuts and cashews are used to make 40 bags of a mixture that sells for $1.15 per bag. How much is a bag of cashews worth
Answer by richwmiller(17219)

p+c=40,
p=c,
P=C-.20,
20*(C-.20)+20*C=40*1.15
Cashews=$1.25
Peanuts=$1.05